# New Paradigm of Adversarial Training: Breaking Inherent Trade-Off between Accuracy and Robustness via Dummy Classes

### v1

@@ C.2 Additional Ablation Study
Hyper-parameter $\alpha$ is designed to adjust the weights of benign and adversarial losses for the model update procedure, thus directly injecting user preference on either clean accuracy or adversarial robustness. In this work, we do not suggest fine-turning $\alpha$ as an approach to impact the trade-off, because our idea implies equal importance of benign and adversarial samples respectively regarding the learning of accuracy and robustness, and our soft label construction has already considered an appropriate balance between them. However, as some previous works on the trade-off problem adopt a similar hyper-parameter (e.g., the regularization parameter $\lambda$ in TRADES), we still provide such an option in case our user does have any specific requirements on it in their practices. Predictably, as illustrated in Figure 9, as $\alpha$ increases, which means more importance is attached to the benign samples, there is overall an increasing (or decreasing) trend for the clean accuracy (or adversarial robustness) of the model. Nevertheless, it is also worth noting that $\alpha=0.5$ as our default one, which means no specific preference on accuracy or robustness, is basically the optimal choice, just as we expected.
Appendix D Limitations
@@
Despite the significant advancement achieved by this work, here we identify two limitations respectively for our new AT paradigm and the proposed DUCAT method, to, hopefully, facilitate future works in this area. Firstly, although the conventional trade-off between clean accuracy and robustness in the current AT paradigm has been broken by our new AT paradigm, with both of them achieving new SOTA performance, we can still observe certain tension between these two objectives, though much slighter than the conventional one. It can be found in the ablation studies illustrated by Figures 7 and 9 that the trend of clean and robust curves have still not fully aligned with each other in certain ranges. This implies that our new AT paradigm does still not harmonize them perfectly, thus just serving as a stepping stone instead of an end-all solution. Secondly, integrating DUCAT with different benchmarks shows different degrees of advancement. Particularly, compared with the remarkable results with PGD-AT and Consistency-AT, as well as the considerable one with MART, the outcomes on TRADES seem less impressive. This is probably because TRADES explicitly balances accuracy and robustness, as well as adopts a unique KL-divergence based regularization term to generate training adversarial samples, which may not fit well with DUCAT. But still, this implies that, despite the effectiveness and simplicity of the proposed DUCAT, there should be room for better specific methods under our new AT paradigm outperforming it, especially for the consistency of effectiveness across different AT benchmarks.

### v2

@@ Appendix F Ablation Study
Ablation Study on $\alpha$ (Figure 13(d)). The hyper-parameter $\alpha$ is designed to explicitly adjust the weights of benign and adversarial losses for the model update procedure, directly injecting the learning preference to attach more importance to either benign or adversarial samples, which is a general strategy not specific to our method and has a predictable impact (i.e., either better accuracy or robustness). In this work, we would not suggest fine-tuning $\alpha$ as an approach to impact the trade-off, because our idea implies equal importance of benign and adversarial samples respectively regarding the learning of accuracy and robustness, and our unique two-hot soft label construction has already considered an appropriate balance between them. However, as some previous works on the trade-off problem adopt a similar hyper-parameter (e.g., the regularization parameter $\lambda$ in TRADES), we still provide such an option in case our user does have any specific requirements to customize it in their practices. Predictably, as illustrated in Figure 13(d), as $\alpha$ increases, which means more importance is attached to the benign samples, there is overall an increasing (or decreasing) trend for the clean accuracy (or robustness) of the model. Nevertheless, it is also worth noting that $\alpha=0.5$ as our default one, which means no specific preference on accuracy or robustness, is basically the optimal choice, just as we expected.
Appendix G Further Discussion
In this section, we provide additional discussion on three aspects where further clarifications might be found useful, namely 1) the difference with DuRM (Wang et al., 2023a), which involves dummy classes in a different way, 2) the comparison with the conventional soft-label strategy, as we propose a novel two-hot soft label, and 3) the specific analysis about TRADES + DUCAT, since the improvements there is not as impressive as others.
G.1 Difference with DuRM
  • Significance. DuRM tends to adopt a smaller $C_{d}$ (e.g., typically 1 or 2) than $C$, and expects there should be no samples actually classified to the dummy classes. In contrast, our paradigm always adopts another $C$ dummy classes, clearly encouraging adversarial samples to be classified into them, and utilizing the one-to-one correspondences between original and dummy classes to detect and recover adversarial samples at inference time, thus acquiring certain robustness.
  • Achievement. Despite AT has been mentioned as one of several application scenarios of DuRM, as it actually does not involve any special designs for AT, the reported improvement is trivial (i.e., no significant improvement compared with PGD-AT). On the contrary, our approach achieves SOTA performance as shown in Section 3.
G.2 Difference between Our Two-Hot Soft Label and Conventional Ones
Careful readers may notice that the suggested two-hot soft label is distinguished from the conventional soft label, such as the ones proposed for label smoothing (Müller et al., 2019; Shafahi et al., 2019; Wu et al., 2024), in both motivation and specific implementation. In short, previous soft label technologies convert one-hot label vectors into one-warm vectors that represent a low-confidence classification (Shafahi et al., 2019). In contrast, the two-hot soft label for DUCAT is neither built in a one-warm format nor aiming at a low-confidence classification. More specifically, from the perspective of motivation, our two-hot soft label is to explicitly bridge corresponding original and dummy classes as the suboptimal alternative target of each other, so that their separation also becomes learnable, while the conventional ones aim to promote learning from the soft target as better guidance that reflects the underlying distribution of data (Müller et al., 2019; Shafahi et al., 2019; Wu et al., 2024). On the other hand, from the perspective of the specific implementation, different from conventional soft label combining one-hot target with uniform or crafted distribution (Müller et al., 2019; Shafahi et al., 2019; Wu et al., 2024), our two-hot soft label just combines two one-hot targets namely as the primary and alternative targets. All in all, the unique two-hot soft label is proposed for the first time, strongly supporting the outstanding effectiveness of DUCAT.
G.3 Analysis about Less Improvement of TRADES + DUCAT than Others
Although DUCAT as an independent AT method (referred to as PGD-AT + DUCAT) achieves SOTA performance, it helps relatively less in enhancing TRADES than other AT benchmarks when serving as a plug (i.e., the improvement of TRADES + DUCAT over TRADES is not impressive as others). In our opinion, this is due to the particular training adversarial samples used by TRADES, which is less appropriate for building dummy clusters than typical ones.
Specifically, different from most AT methods including the other three benchmarks, PGD-AT, MART and Cons-AT, that by default generate training adversarial samples by PGD-10, TRADES particularly crafts training adversarial samples through maximizing its own KL-divergence regularization term (Zhang et al., 2019). Because TRADES defines a boundary error measured through this KL term, which occurs when specific data points are sufficiently close to the decision boundary that can easily cross it under slight perturbation (Zhang et al., 2019), thus it makes sense to accordingly craft training adversarial samples to reduce this error. But as a consequence, these training adversarial samples naturally focus more on boundary establishment instead of reasonable data distribution patterns, which predictably degrades our DUCAT, as DUCAT assumes most of the adversarial samples from one class should belong to the corresponding dummy class.
Appendix H Limitations
Despite the significant advancement achieved by this work, here we identify two limitations respectively for our new AT paradigm and the proposed DUCAT method. We hope these discussions could facilitate future works on the trade-off problem between clean accuracy and adversarial robustness among the AT community.
@@ Appendix H Limitations
Firstly, although the conventional trade-off between accuracy and robustness in the current AT paradigm has been released by our new AT paradigm, with both of them achieving new SOTA performance, we can still observe certain tension between these two objectives, though much slighter than before. It can be found in the ablation studies illustrated by Figures 13 that the trend of clean and robust curves has still not fully aligned with each other in certain ranges. This implies that our new AT paradigm still does not harmonize them perfectly, thus just serving as a stepping stone instead of an end-all solution. Also, while the clean accuracy under DUCAT is significantly improved compared with the current AT methods, there is still a gap compared with standard training (e.g., the SOTA accuracy of ResNet-18 on CIFAR-10 in the clean context is about 95%).
Secondly, integrating DUCAT with different benchmarks shows different degrees of advancement. As discussed in Appendix G.3, compared with the remarkable results with PGD-AT and Cons-AT, as well as the considerable one with MART, the outcomes on TRADES seem less impressive. This implies that, despite the effectiveness and simplicity of the proposed DUCAT, there should be room for better specific methods under our new AT paradigm to outperform it, especially regarding the different mechanisms of specific benchmarks like TRADES.